Chorizo:

Wesley is making a patio from stones of two sizes, 5 inches wide and 10 inches wide. He wants to begin and end his pattern with a 10-inch stone so there will be one more of the 10-inch stones than of the 5-inch stones. His patio will be 145 inches wide. How many 10-inch stones will Wesley need for one row?

Timmyspu:

Welcome to question cove I hope you enjoy your time here if you have any questions feel free to ask. So what information do we already know?

Walen:

do u have a picture of the question? because i dont really understand it by a patio do u mean more than one row of stones or just one row of stones

Walen:

by wide do u mean cm^2 or just cm

surjithayer:

10+5=15 145-10=135 135/15=9 so there will be 9 stones of 5 inches length and 9+1=10 stones of 10 inches length.